where, W = Wheat yield (kg/ha), F = Fertiliser (kg/ha) and L = Labour (hours/ha) 

Derive an expression for the least-cost expansion path given that fertiliser costs R12 per kg and labour costs R6 per hour. [4]

If the wheat producer has a budget of R9 000 to spend on fertiliser and labour, how much of each input should he use and what is the quantity of wheat he will produce? [4]

Derive expressions for the value of marginal product (VMP) of fertiliser and labour given that wheat sells at R72 per kg.


1
Expert's answer
2021-09-09T17:14:24-0400

"\frac{∂W/∂L}{∂W/∂F}=\frac{P_L}{P_F} \\

or \; \frac{MP_L}{MP_F}= \frac{P_L}{P_F} \\

\frac{F^{1/2}L^{-3/4}}{2F^{-1/2}L^{1/4}} = \frac{6}{12} \\

\frac{L^*}{F^*} = 1 \\

or \; L^*=F^*"

which is the expression for the least cost expansion path

Now,

"C=P_fF + P_lL \\

C=12F+6L \\

C=9000 \\

L^*=F^* \\

18L^*=9000 \\

L^* = 500 \;units \\

F^*=500 \;units"

The total production

"W^* = 4F^{*1/2}L^{*3/4} \\

=4(500)^{3/4} \\

= 422.95 \;units"

which is the optimum quantity produced

Now, "VMP =P_W \times MP_L"

"MP_L= F^{1/2}L^{-3/4} \\

P_W = R72 \;per \;kg \\

VMP = P_W \times MP_L \\

= \frac{72F^{1/2}}{L^{3/4}}"